Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

An urn contains 10 balls, numbered from 1 to 10. If 2 balls

Expert replies
by BTGmoderatorDC » Sat Jul 20, 2019 4:17 pm
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

An urn contains 10 balls, numbered from 1 to 10. If 2 balls are selected at random with replacement from the urn, what is the probability that the sum of the 2 numbers on the balls will be even?

A. 25%
B. 37.5%
C. 50%
D. 62.5%
E. 75%

OA C

Source: Veritas Prep
Join the discussion
Source: — Problem Solving |

by Darkknightreturning » Sat Jul 20, 2019 9:18 pm
For the sum to be even, the balls selected must be both odd or both even.
Both odd numbered balls will mean, balls with numbers 1,3,5,7,9 i.e. 5
Both even numbered balls will mean balls with numbers 2,4,6,8,10 i.e. 5

Any other combination will yield an odd sum.
For selecting 2 odd numbered balls, number of ways will be 5C2
For selecting 2 even numbered balls, number of ways will be 5C2
Total ways of selecting 2 balls out of 10 balls = 10C2

Thus, probability of getting balls with even sum = (5C2+5C2)/10C2 which works out to be 44.44%.
Join the discussion

by Darkknightreturning » Sat Jul 20, 2019 9:23 pm
BTGmoderatorDC wrote:An urn contains 10 balls, numbered from 1 to 10. If 2 balls are selected at random with replacement from the urn, what is the probability that the sum of the 2 numbers on the balls will be even?

A. 25%
B. 37.5%
C. 50%
D. 62.5%
E. 75%

OA C

Source: Veritas Prep
Ok. since the question includes the term "with replacement". Thus, once one ball has been taken out, it also gets replaced.
So the number of ways in which odd numbered balls are selected = 5x5 (since the ball gets replaced)
the number of ways in which even numbered balls are selected = 5x5 (since the ball gets replaced)

Total ways of selecting 2 balls out of 10 balls with replacement = 10x10
Thus, the probability of getting balls with even sum = (25+25)/100= 50%
Thus, C is the correct answer.
Join the discussion

by Brent@GMATPrepNow » Sun Jul 21, 2019 4:53 am
BTGmoderatorDC wrote:An urn contains 10 balls, numbered from 1 to 10. If 2 balls are selected at random with replacement from the urn, what is the probability that the sum of the 2 numbers on the balls will be even?

A. 25%
B. 37.5%
C. 50%
D. 62.5%
E. 75%

OA C

Source: Veritas Prep
There are 2 possible ways to get an EVEN sum:
1) The 1st and 2nd balls are both EVEN
2) The 1st and 2nd balls are both ODD

So, P(sum is even) = P(1st is even AND 2nd is even OR 1st is odd AND 2nd is odd)
= P(1st is even AND 2nd is even) + P(1st is odd AND 2nd is odd)
= [P(1st is even) x P(2nd is even)] + [P(1st is odd) x P(2nd is odd)]
= [1/2 x 1/2] + [1/2 x 1/2]
= 1/4 + 1/4
= 1/2
= 0.5
= 50%

Answer: C

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by Scott@TargetTestPrep » Tue Jul 30, 2019 8:38 am
BTGmoderatorDC wrote:An urn contains 10 balls, numbered from 1 to 10. If 2 balls are selected at random with replacement from the urn, what is the probability that the sum of the 2 numbers on the balls will be even?

A. 25%
B. 37.5%
C. 50%
D. 62.5%
E. 75%

OA C

Source: Veritas Prep

The sum of the 2 numbers will be even if the 2 numbers chosen are both even or both odd. The probability of choosing 2 even numbers is 1/2 x 1/2 = 1/4. Likewise, the probability of choosing 2 odd numbers is 1/2 x 1/2 = 1/4. Therefore, the probability of choosing 2 numbers whose sum is even is 1/4 + 1/4 = 1/2.

Answer: C

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion